jhonadam1:

The sum of three consecutive even integers is –78. What are the integers? Describe how you would use the strategy of writing an equation to solve the given problem.

Narad:

Do you know what is an even integer?

surjithayer:

let the consecutive even integers be x,x+2,x+4 x+x+2+x+4=-78 3x=-78-6 3x=-84 x=? x+2=? x+4=?